OpenStudy (anonymous):

2.93 × 10^−7 in decimal form?

OpenStudy (mathstudent55):

A negative exponent of 10 means you move the decimal point that many spaces to the left. For example: 4.98 * 10^(-5) means take 4.98, and move the decimal point 5 places to the left. 4.98 * 10^(-5) = 0.0000498 Now try it with your problem.

OpenStudy (anonymous):

What is the sum written in scientific notation?\[2.9x10^{10}+5.35x10^{10}\]

OpenStudy (mathstudent55):

Just add the first parts of the two numbers. Then use the same power of 10. For example: \(2.5 \times 10^7 + 3.8 \times 10^7 = (2.5 + 3.8) \times 10^7 = 6.3 \times 10^7\)